### v1

# MAUGIF: Mechanism-Aware Unsupervised General Image Fusion via Dual Cross-Image Autoencoders

Abstract
Image fusion aims to integrate structural and complementary information from multi-source images. However, existing fusion methods are often either highly task-specific, or general frameworks that apply uniform strategies across diverse tasks, ignoring their distinct fusion mechanisms. To address this issue, we propose a mechanism-aware unsupervised general image fusion (MAUGIF) method based on dual cross-image autoencoders. Initially, we introduce a classification of additive and multiplicative fusion according to the inherent mechanisms of different fusion tasks. Then, dual encoders map source images into a shared latent space, capturing common content while isolating modality-specific details. During the decoding phase, dual decoders act as feature injectors, selectively reintegrating the unique characteristics of each modality into the shared content for reconstruction. The modality-specific features are injected into the source image in the fusion process, generating the fused image that integrates information from both modalities. The architecture of decoders varies according to their fusion mechanisms, enhancing both performance and interpretability. Extensive experiments are conducted on diverse fusion tasks to validate the effectiveness and generalization ability of our method. The code is available at https://anonymous.4open.science/r/MAUGIF.
1 Introduction
Image fusion aims to integrate complementary information from multiple source images into a single output that is more informative, visually interpretable, and suitable for downstream tasks such as object detection [He2023], segmentation [Wang2025DiFusionSegDS], and remote sensing analysis [Hong2018]. General image fusion refers to designing a unified model or framework that can be applied to various types of image fusion tasks. These tasks include hyperspectral-multispectral image fusion (HMF), visible-infrared image fusion (VIF), multi-focus image fusion (MFF), and medical image fusion (MEF). The HMF task seeks to fuse a low-spatial-resolution hyperspectral image (HSI) with a high-spatial-resolution multispectral image (MSI) to produce a high-spatial-resolution HSI [Hysure]. The VIF integrates thermal radiation data from infrared sensors with rich texture and color information from visible cameras. This results in the fused images that simultaneously highlight salient objects and retain visual fidelity [Ma2021GANMcCAG]. The MFF task aims at generating an all-in-focus image from a series of partially focused images [Quan2025]. The purpose of MEF is to display the fine structure and functional metabolism of the lesion simultaneously, thereby assisting doctors in making more accurate diagnoses and assessment [HE2025102666].
Figure 1: The number of model parameters and inference time of different fusion methods.
Existing image fusion methods can be broadly categorized into task-specific and general-purpose. Task-specific methods are specially designed for a particular type of image fusion task. These methods typically leverage prior image information such as sparsity [OTD] and structural sparsity [S4], or deep image features [ZSL] to enhance the fusion performance [Zhang2023]. However, these methods often exhibit limited generalization capability across different fusion tasks. In contrast, general image fusion methods aim to adopt a uniform fusion strategy across tasks [Xu2020U2FusionAU]. These methods generally use general features including frequency coefficients [Wang2024a] and deep features [Liang2024FusionFD] to optimize the model. Nevertheless, most existing methods fail to fully account for the underlying mechanisms of different tasks, thereby limiting their adaptability and interpretability [Zhu2024TaskCustomizedMO].
To address these issues, we propose a mechanism-aware unsupervised general image fusion (MAUGIF) method based on Dual Cross-Image Autoencoders (DCIAE). According to the inherent mechanisms of different fusion tasks, we introduce a classification of additive and multiplicative fusion. For example, the fused output of VIF can be modeled as a combination where thermal saliency complements visible details in an approximately additive manner, thus belonging to additive fusion. In contrast, HMF involves spatial deblurring and super-resolution of HSIs, which can be modeled through multiplicative interactions rather than direct addition. Then, dual encoders are employed to extract the common content from source images, while modality-specific details are isolated. The dual decoders then act as feature injectors, reconstructing images by reintegrating their unique characteristics into the shared content. In the fusion process, the modality-specific features are injected into the source image, thereby generating the fused image that integrates information from both modalities. Notably, different fusion mechanisms correspond to distinct decoder architectures, i.e., different feature injection manners, enhancing both the performance and interpretability. Moreover, our model consists of two lightweight encoder–decoder branches, ensuring a compact architecture. The fusion process is accomplished in a single decoding step, which eliminates the need for designing complex fusion rules. As demonstrated in Figure 1, this design leads to high computational efficiency. This makes our method well-suited for real-time applications and deployment in resource-constrained environments. We summarize the main contributions of this paper as follows.
  • An interpretable and efficient general image fusion framework. We propose a mechanism-aware general image fusion method termed MAUGIF, which explicitly disentangles the shared content and modality-specific details of the source images. As shown in Figure 2, MAUGIF enables direct visualization of the fusion process, and the contribution of each source image can be clearly observed. In addition, our model is lightweight and the fused image is generated in a single decoding step, leading to high computational efficiency.
  • A mechanism-aware taxonomy of fusion tasks. By proposing a classification of fusion into additive and multiplicative types, we can gain insight into how the modality-specific features are integrated. This not only guides the architectural design of the network, but also enhances model generalization and interpretability.
  • Comprehensive validation across diverse fusion tasks. Extensive numerical experiments are conducted on various fusion tasks (HMF, VIF, MFF and MEF) to validate the effectiveness and generalization capability of the proposed general image fusion method.
Figure 2: Visual illustration of the fusion process for MFF.
Figure 3: The mechanism-aware DCIAE framework for general image fusion, illustrated with four scenarios.
2 Related work
2.1 Task-specific image fusion methods
Task-specific fusion methods are designed for a particular fusion task, leveraging the inherent properties of image and priors to enhance performance. Model-based fusion methods typically rely on mathematical modeling and optimization to exploit the priors of images [NSSR, OTD]. For example, these methods often leverage various matrix or tensor decompositions to capture and exploit the intrinsic structure of image data [CP1, CSTF]. In addition, they can incorporate image priors via designed regularizers, such as low-rankness [Hysure], sparsity [WANG2014477], smoothness [Ma2016], and non-local self-similarity [LTMR]. Deep learning-based methods capture high-level semantic content by learning hierarchical representations directly from data. Numerous methods employ convolutional neural networks (CNNs) to extract such features from source images [ZSL, Li2019]. Subsequently, generative adversarial networks (GANs) [Ma2019FusionGANAG, Zhang2021a] and vision Transformers [Shao2024, Liu2024] have been introduced to better preserve structural details and global contextual information. More recently, diffusion-based fusion methods have also attracted significant attention [Yue2023, Wu2025]. Furthermore, high-level vision tasks including object detection and semantic segmentation, have been incorporated to guide the fusion process [Tang2022ImageFI, Wang2025DiFusionSegDS]. These methods often deliver optimal performance on specific fusion tasks but exhibit limited generalizability.
2.2 General image fusion methods
General image fusion aims to handle multiple fusion tasks within a unified framework [Ma2022SwinFusionCL]. Recently, Zhang et al. [Zhang2020a] proposed a supervised training framework for MFF, which can be generalized to other tasks by adjusting fusion conditions. Ma et al. [Ma2022] employed a unified network architecture and loss function, but trained separate models for each individual task. Wang et al. [Wang2024a] proposed a semi-supervised learning approach, which decomposes the source images into features of different frequencies and adaptively merges the features. The work in [Zhu2024] accommodated diverse fusion tasks by dynamically customizing a mixture of task-specific adapters, preserving task individuality while leveraging cross-task commonalities.
3 The proposed method
In this paper, the source images are denoted as X and Y, respectively, and the fused image is denoted as F.
3.1 Mechanism-Aware Image Fusion
To capture the diverse fusion mechanisms of information integration, we classify image fusion into additive and multiplicative types. The classification depends on whether the injected modality-specific features are added to the target or modulate it. The specific definitions are as follows.
Definition 3.1
(Additive fusion) Given source images X and Y capturing the same scene, additive fusion aims to generate a composite image F such that:
$\vskip-2.84544pt\textbf{F}=\textbf{Y}+\psi(\Delta_{x})\quad\text{or}\quad\textbf{F}=\textbf{X}+\psi(\Delta_{y}),$ (1)
where $\psi$ is a nonexpansive projection, i.e., $\|\psi(x)-\psi(y)\|\leq\|x-y\|$, $\Delta_{x}:=\textbf{X}-\textbf{C}$ and $\Delta_{y}:=\textbf{Y}-\textbf{C}$ are the modality-specific features unique to X and Y, respectively, and C is the common content of the same size as X and Y.
For example, in visible-infrared image fusion, the goal is to highlight the thermal radiation of the target from the infrared image (IR) while preserving the structural content of the visible image, which is typically considered additive fusion. However, IR image often contains redundant regions and noise, which should not be blindly injected into the fused image. Therefore, $\psi$ acts as a thresholding mechanism to suppress irrelevant responses in modality-specific components. The non-expansiveness of $\psi$ ensures that the features are selectively preserved, but never amplified.
Remark 3.1
(The choice of $\psi$) The design of $\psi$ is adapted to the specific fusion task. For MFF where the modality-specific information of images is all needed, $\psi$ can be set to the identity mapping, i.e. $\psi(x)=x$, to preserve all fine details. For VIF, a hard-thresholding:
$\psi(x)=x\quad\text{if}~~x>\sigma;\quad\text{otherwise}~~0,\vskip-2.84544pt$ (2)
can be selected to suppress weak or spurious responses and enhance salient thermal targets, where $\sigma\in[0,~0.4]$.
Definition 3.2
(Multiplicative fusion) Given source images X and Y capturing the same scene, multiplicative fusion generates the fused image F via modulation like:
$\textbf{X}=\textbf{R}*\textbf{F}\quad\text{and}\quad\textbf{Y}=\textbf{F}*\textbf{B},\vskip-2.84544pt$ (3)
where R and B are degradation fields, ‘*’ denotes matrix multiplication or Hadamard product, etc. Equation (3) indicates that X and Y are different degraded versions of F.
For instance, hyperspectral-multispectral image fusion (HMF) model is typically formulated as:
$\vskip-2.84544pt\min_{\textbf{F}}\|\textbf{FB}-\textbf{X}\|_{F}^{2}+\|\textbf{RF}-\textbf{Y}\|_{F}^{2},\vskip-2.84544pt$
where X and Y denote the HSI and MSI, B and R represent the spatial and spectral degradation operators, respectively [Hysure]. Clearly, HMF belongs to multiplicative fusion.
3.2 Dual cross-image autoencoders framework
To achieve general image fusion, it is crucial to identify the fundamental commonalities shared across diverse image fusion tasks. We observe that:
  • Due to modality differences, images X and Y are inherently distinct, each with unique features, i.e., $\textbf{X}\neq\textbf{Y}$.
  • Despite modality differences, $\mathbf{X}$ and $\mathbf{Y}$ correspond to the same scene and thus share common information, which we denote as $\mathcal{C}(\mathbf{X},\mathbf{Y})\neq\emptyset$, where $\mathcal{C}(\cdot,\cdot)$ represents the set of shared semantic or structural components.
Based on this, we propose the dual cross-image autoencoder (DCIAE) framework for general image fusion, as displayed in Figure 3. The process can be divided into three steps:
1) To extract the common components of source images, we employ two encoders, denoted as $E_{x}(\cdot)$ and $E_{y}(\cdot)$, to encode X and Y respectively, such that $E_{x}(\textbf{X})\approx E_{y}(\textbf{Y})$. The loss function for the encoders can be formulated as:
$loss_{e}=loss_{1}+\lambda loss_{2},$ (4)
where $\lambda>0$ is a weight parameter,
$\vskip-3.0ptloss_{1}:=\|E_{x}(\textbf{X})-E_{y}(\textbf{Y})\|_{F}^{2},$ (5)
$loss_{2}:=\|E_{x}(\textbf{X})-\textbf{X}\|_{F}^{2}+\|\text{E}_{y}(\textbf{Y})-\textbf{Y}\|_{F}^{2}.$ (6)
Here, $loss_{1}$ encourages the encoders to extract the shared content between images $\mathbf{X}$ and $\mathbf{Y}$, while $loss_{2}$ prevents degenerate solutions like $E_{x}(\mathbf{X})=E_{y}(\mathbf{Y})=0$. Notably, $loss_{2}$ ensures that the learned common components remain faithful to the source images and enforces it to have the same size as the source images. This aims primarily to extract common content that aligns with visual intuition.
2) Two decoders, denoted as $D_{x}(\cdot)$ and $D_{y}(\cdot)$, are employed to decode the learned shared content. The objective is to reconstruct the source images X and Y from the common components, i.e.,
$D_{x}(E_{x}(\textbf{X}))=\textbf{X},\quad D_{y}(E_{y}(\textbf{Y}))=\textbf{Y}.$ (7)
The decoder acts as a ‘feature injector’, which injects the source image’s modality-specific features into the shared content, thereby reconstructing the source images. The loss function for the decoders can be defined as:
$loss_{d}:=\|D_{x}(E_{x}(\textbf{X}))-\textbf{X}\|_{F}^{2}+\|D_{y}(E_{y}(\textbf{Y}))-\textbf{Y}\|_{F}^{2}.$ (8)
Their architecture is shown in Figure 4.
Figure 4: Architecture of the Encoder and Decoder.
3) Finally, we perform image fusion by injecting the modality-specific feature of image X into image Y, or vice versa. Therefore, the fused image F can be formulated as:
$\textbf{F}=D_{x}(\textbf{Y})\quad\text{or}\quad\textbf{F}=D_{y}(\textbf{X}).$ (9)
In this process, the decoder serves directly as a fusion operator, shifting the fusion process from conventional rule-based design to intrinsic feature injection.
Additive fusion and multiplicative fusion correspond to different decoder architectures, as shown in Figure 3 (e). In additive fusion, the decoder architecture follows a residual paradigm, where the modality-specific features of X are injected as an additive manner, i.e.,
$\vskip-1.42271ptD_{x}(\cdot)=(\cdot)+\psi(\Delta_{x}).$ (10)
In multiplicative fusion, the decoder instead generates a modulation field to refine the structure of the input image.
Remark 3.2
Under DCIAE framework, image fusion is no longer a ‘black-box’ process. We can clearly observe the modality-specific features contributed by the source images during the fusion process. Moreover, DCIAE framework is few shot, requiring only one image pair for model training. In addition, the fusion process involves just a single decoding step, making it highly efficient.
Next, we analyze the factors affecting DCIAE’s performance. Taking additive fusion as an example, we denote
  • $\mathcal{C}^{*}$: the ground-truth shared content between X and Y.
  • $\widehat{\textbf{F}}$: the fusion result of DCIAE under ideal decoder, i.e., $D_{x}(E_{x}(\textbf{X}))=\textbf{X}$ holds exactly.
  • $\textbf{F}^{*}$: The optimal fusion result under DCIAE framework.
Assume that $\psi$ is a truncation function selected as needed. Then, by combing (9) with (10), one has:
F $=\textbf{Y}+\psi(D_{x}(E_{x}(\textbf{X}))-E_{x}(\textbf{X})),$
$\widehat{\textbf{F}}$ $=\textbf{Y}+\psi(\textbf{X}-E_{x}(\textbf{X})),\quad\textbf{F}^{*}=\textbf{Y}+\psi(\textbf{X}-\mathcal{C}^{*}).$
According to the non-expansiveness of $\psi$, one has
$\|\textbf{F}-\textbf{F}^{*}\|$ $\leq\|\textbf{F}-\widehat{\textbf{F}}\|+\|\widehat{\textbf{F}}-\textbf{F}^{*}\|$
$\leq\|D_{x}(E_{x}(\textbf{X}))-\textbf{X}\|+\|E_{x}(\textbf{X})-\mathcal{C}^{*}\|.$
This indicates that the performance of the DCIAE framework is mainly influenced by two factors: the network’s expressive capacity and the accuracy of the shared content.
4 Experiments
In this section, we validate the effectiveness of the proposed MAUGIF method on four tasks, including hyperspectral-multispectral image fusion (HMF), visible-infrared image fusion (VIF), multifocus image fusion (MFF), and medical image fusion (MEF). We also conduct a comparison of computational complexity and ablation experiments.
4.1 Experimental setup
4.1.1 Implementation details
All experiments are conducted on Ubuntu 20.04 with the PyTorch 2.2.0. The hardware configuration includes an Intel Core i7-13700KF CPU with a clock speed of 3.4GHz, 32GB of RAM, and an NVIDIA GeForce RTX 4090 GPU with 24GB of VRAM. Throughout the training process, the AdamW [Loshchilov2017DecoupledWD] optimizer is employed, along with a linear learning rate schedule that incorporates a warm-up phase. The maximum learning rate is set to $1\times 10^{-4}$, the batch size is set to 32, and the training runs for 50 epochs.
4.1.2 Dataset
Next, we introduce the datasets used in our experiments.
CAVE1 and Harvard Dataset [data11] for HMF: For the CAVE dataset, we selected ‘Toys’ and ‘Peppers’, whose spatial dimensions were downsampled by different scale factors (sf). We simulated the MSIs with three bands by the Nikon D700 camera2. For Harvard dataset, the ‘img3’ and ‘imgc6’ images are selected. Each contains a 31-band HSI, which serves as the ground truth. Then, the low-resolution HSIs are generated through downsampling the original HSI at different scale factors. We divide the spectral bands of HSI into three parts and generate a 3-band MSI by averaging the bands. We corrupt the HSI and MSI with Gaussian noise to obtain SNRs of 35 dB and 40 dB, respectively.
MSRS Dataset [Tang2022ImageFI] for VIF: This dataset is a benchmark for infrared-visible image fusion and segmentation. It comprises 1,444 well-registered image pairs. Following the official protocol, we adopt the original split with 1,083 pairs for training and 361 pairs for testing. All images have a resolution of 640×480 and are accompanied by pixel-level semantic segmentation annotations.
MFI-WHU Dataset [zhang2021mff] for MFF: This dataset is a benchmark for multi-focus image fusion, which contains 120 pairs of images, synthesized with Gaussian blurring and provided with manually annotated decision maps.
Harvard Dataset3 for MEF: This dataset is a well-known medical imaging benchmark for multimodal fusion. It provides a large number of CT–MRI, PET–MRI, and SPECT–MRI image pairs, which are widely used for training and evaluating medical image fusion models.
11footnotetext: https://www.cs.columbia.edu/CAVE/databases/multispectral/ 22footnotetext: https://maxmax.com/spectral_response.htm33footnotetext: http://www.med.harvard.eduAANLIB/home.html
((a)) HSI/MSI
((b)) CNMF
((c)) Hysure
((d)) ZSL
((e)) TriDLR
((f)) DTDNML
((g)) MAUGIF
((h)) GT
Figure 5: The fused images of compared methods and the corresponding error maps for ‘Toys’ dataset.
Figure 6: Qualitative comparison of MAUGIF with 9 SOTA methods on image 00500D from the MSRS visible-infrared dataset.
4.2 Comparison on HMF task
In hyperspectral-multispectral image fusion (HMF), the MAUGIF approach performs fusion without relying on the knowledge of the degradation process, and thus falls into the category of blind fusion [CP1]. Therefore, we select five blind HMF methods for comparison, including the classical CNMF [CNMF] and HySure [Hysure], as well as the recently proposed ZSL [ZSL], TriDLR [Yang2025], and DTDNML [Wang2024].
Table 1: The evaluation metrics of the fused images for HMF.
Dataset Method PSNR SSIM ERGAS SAM Time(s)
CNMF 32.03 0.881 2.199 17.69 9.9
Hysure 28.43 0.866 3.096 23.74 40.5
Peppers ZSL 40.32 0.978 1.172 11.95 387.1
 (sf=16) TriDLR 36.85 0.952 1.353 15.19 57.3
DTDNML 37.42 0.969 1.302 14.79 31.8
MAUGIF 40.91 0.982 1.038 11.87 10.7
CNMF 29.75 0.872 0.647 16.79 8.6
Hysure 27.01 0.832 0.779 21.74 36.1
 Toys ZSL 38.27 0.964 0.408 11.37 325.8
(sf=32) TriDLR 34.50 0.948 0.438 12.43 46.9
DTDNML 34.86 0.953 0.436 13.28 27.5
MAUGIF 39.68 0.980 0.297 9.22 8.9
CNMF 38.59 0.943 3.734 11.97 11.2
Hysure 40.161 0.949 3.512 8.69 56.5
 img3 ZSL 42.32 0.972 2.528 8.75 487.5
(sf=8) TriDLR 41.97 0.969 2.803 9.48 71.4
DTDNML 41.06 0.970 2.742 9.66 38.7
MAUGIF 42.90 0.975 2.431 9.07 12.4
CNMF 37.14 0.954 0.923 6.02 9.4
Hysure 37.87 0.957 0.983 5.79 48.3
 imgc6 ZSL 39.82 0.967 0.867 5.45 420.3
(sf=16) TriDLR 38.64 0.962 0.927 4.92 64.7
DTDNML 37.76 0.958 0.948 5.85 34.9
MAUGIF 40.76 0.969 0.853 5.03 10.8
Table 1 presents the fusion performance of the compared methods on four HSI datasets. From the table, the CNMF and HySure exhibit relatively weaker performance, particularly under large scale factor cases. In this cases, the fusion model becomes severely ill-posed, which limits their performance. TriDLR reduces the number of parameters through tensor triple decomposition. It adopts a strategy of estimating the degradation operator and performing fusion simultaneously, effectively avoiding error accumulation and achieving superior performance. The ZSL achieves a relatively accurate estimation of the degradation operators, which results in less error accumulation during its subsequent fusion process and leads to competitive fusion performance. However, when the scale factor (sf) is large, the accuracy of the degradation operator estimated by ZSL decreases, leading to a significant drop in fusion performance. In contrast, the proposed MAUGIF method exhibits greater robustness to large sf, as the proposed method can perform spectral super-resolution on the MSI, which is less sensitive to spatial downsampling. Overall, MAUGIF achieves the optimal performance in four datasets, which demonstrates its effectiveness in HMF.
Figure 7: Experimental results of compared methods on three multi-focus image datasets and the corresponding PSNR values.
((a))
((b))
((c))
Figure 8: Experimental results of compared methods on three types of medical image datasets. (a)–(c) correspond to different modalities.
Figure 5 presents the fused images and error maps on the ‘Toys’ dataset for visual comparison. As shown in the figure, the fused images of HySure exhibit noticeable color distortions, which indicates a significant estimation error in the spectral degradation operator. As can be seen from the error maps, ZSL achieves competitive results. However, MAUGIF exhibits the fewest artifacts and color distortion, achieving the best visual reconstruction quality.
4.3 Comparison on VIF task
To evaluate the performance of the MAUGIF on visible-infrared image fusion task, we conduct experiments on MSRS [Tang2022ImageFI] dataset, and compare the MAUGIF with several state-of-the-art (SOTA) fusion methods. The quantitative results presented in Figure 2 demonstrate that MAUGIF achieves five top rankings and one second-place ranking among all competing methods. Specifically, it records an MI of 5.638, an SCD of 1.731, a VIF of 0.949, a $Q^{AB/F}$ score of 0.652, and an SSIM of 0.482. These results indicate that MAUGIF effectively preserves mutual information, structural similarity, and visual information fidelity, suggesting superior retention of texture, structure, and perceptual quality.
Table 2: Quantitative comparisons on image pairs from the MFI-WHU and Harvard datasets.
Category MI↑ AG↑ SCD↑ VIF↑ $Q^{AB/F}$ SSIM↑
U2Fusion20 [Xu2020U2FusionAU] 1.988 2.515 1.145 0.516 0.386 0.381
GANMcC21 [Ma2021GANMcCAG] 2.459 2.188 1.458 0.612 0.326 0.309
SDNet21 [Zhang2021SDNetAV] 1.708 2.675 0.986 0.499 0.377 0.364
TarDal22 [Liu2022TargetawareDA] 2.194 1.716 0.697 0.406 0.172 0.238
SegMif22 [Liu2023MultiinteractiveFL] 2.472 3.336 1.594 0.774 0.565 0.339
DDFM23 [Zhao2023DDFMDD] 2.467 2.700 1.556 0.684 0.484 0.327
Dif-Fusion23 [Yue2023DifFusionTH] 3.326 3.889 1.592 0.827 0.583 0.448
DCINN23 [Wang2023AGP] 3.328 3.334 1.478 0.810 0.568 0.369
MRFS24 [Zhang2024MRFSMR] 3.068 3.034 1.288 0.74 0.484 0.350
A2RNet25 [li2025a2rnet] 3.334 2.814 1.481 0.711 0.438 0.359
MAUGIF 5.638 3.457 1.731 0.949 0.652 0.482
Although Dif-Fusion achieves a slightly higher AG of 3.889, MAUGIF still delivers a competitive AG value of 3.457 and consistently outperforms the other methods across all remaining metrics. These observations validate that MAUGIF can generate fused images with enhanced information content, sharper detail representation, and superior perceptual quality compared with existing SOTA approaches.
To better illustrate the fusion performance of MAUGIF, we provide a qualitative comparison, as shown in Figure 6. We can observe that our method preserves fine-grained texture details more effectively than compared approaches. Specifically, in the regions highlighted by the red and green boxes, the text remains sharp, structurally continuous, and clearly recognizable in our results. In contrast, other methods tend to introduce noise interference from the infrared modality, resulting in blurred textures, weakened edge contours, and noticeable degradation of structural fidelity. These visual comparisons further verify the robustness of our fusion strategy in maintaining both clarity and semantic consistency.
In summary, both quantitative and qualitative results demonstrate that MAUGIF preserves complementary information from the source images while maintaining visual fidelity. The fused images retain key details from both visible and infrared modalities, ensuring high visual quality and enhancing the overall integrity of the fusion process.
4.4 The performance on MFF and MEF
Next, we verify the effectiveness and generalization capability of the proposed MAUGIF model on MFF and MEF tasks. Three general image fusion methods are compared, including U2Fusion [Xu2020U2FusionAU], SDNet [Zhang2021SDNetAV], and TC-MOA [Zhu2024TaskCustomizedMO]. The quantitative results are presented in Table 3. From the table, the fused images of MFF obtained by the proposed MAUGIF method significantly outperform the compared methods in terms of both PSNR and SSIM metrics. For example, its average PSNR is 4.5 higher than that of the suboptimal SDNet. For MEF, although there is no ground truth image available as a reference, our method achieves the best overall performance in many metrics. This indicates that the fused image preserves both mutual information and visual information fidelity, which fully validates the effectiveness and generalization of our approach.
Table 3: Quantitative comparisons of six metrics on 361 image pairs from the MSRS dataset.
Method U2Fusion20 SDNet21 TC-MOA24 MAUGIF
Comparison on MFF (with GT reference)
PSNR↑ 28.374 32.153 31.896 36.632
SSIM↑ 0.917 0.955 0.949 0.986
Comparison on MEF (without GT reference)
MI↑ 2.536 2.651 2.818 3.459
AG↑ 5.088 5.265 7.501 7.687
SCD↑ 0.216 0.812 1.235 1.288
VIF↑ 0.411 0.385 0.547 0.563
SF↑ 17.103 23.252 24.276 28.851
Figure 7 and 8 present visual comparisons of the compared methods on two multi-focus image pairs and three types of medical image datasets. For MFF, our method generates fused images with consistently sharp focus across the entire scene, effectively integrating the in-focus regions from all source inputs. Moreover, magnified views of the results reveal that fine details are well preserved with high visual sharpness, whereas other compared methods often suffer from blur or incomplete detail transfer, failing to retain all salient structures effectively. For MEF task, our method successfully integrates complementary information from different brain imaging modalities. Taking MRI–PET image pairs as an example, U2Fusion and TC-MOA introduce noticeable distortions, while U2Fusion and SDNet fail to preserve fine anatomical structures from the MRI modality, particularly in the red regions, where textures appear blurred or weakened. In contrast, our fused results effectively retain both the structural details of MRI and the metabolic activity patterns of PET, demonstrating superior information preservation. This highlights the effectiveness of MAUGIF in MEF task.
4.5 Complexity Discussion
To evaluate the efficiency of the proposed MAUGIF, we compare it against several SOTA models. We measure performance in terms of floating-point operations (FLOPs), model parameters (Params), and per-frame inference time (Time), as summarized in Table 4.
From the table, MAUGIF exhibits a significant advantage in computational efficiency compared with existing fusion networks. For example, MAUGIF contains only 0.00569M parameters and requires 1.76G FLOPs, which is much lower than other methods. In terms of inference latency, MAUGIF achieves an average processing time of 10.63 ms per image pair, making it the fastest among all compared methods. These results demonstrate that MAUGIF achieves a superior balance between fusion quality and computational cost, making it highly suitable for real-time or resource-constrained applications, such as embedded deployment in intelligent perception systems.
Table 4: Quantitative comparison of different methods in terms of FLOPs, Params, Time on the MSRS dataset.
Method Input Size Params/M↓ FLOPs/G↓ Time/ms↓
U2Fusion20[Xu2020U2FusionAU] 480x640 0.66 518.00 695.64
SDNet21[Zhang2021SDNetAV] 480x640 0.07 64.55 122.33
DDFM22[Zhao2023DDFMDD] 480x640 552.81 1.34M 102760
Dif-Fusion23[Yue2023DifFusionTH] 480x640 416.47 1056.86 809.41
SegMiF23[Liu2023MultiinteractiveFL] 480x640 45.63 359.41 309.70
MRFS24[Zhang2024MRFSMR] 480x640 134.96 139.13 124.72
TC-MoA24[Zhu2024TaskCustomizedMO] 480x640 340.35 3145.68 305.35
DiFusionSeg25[Wang2025DiFusionSegDS] 480x640 40.93 174.67 77.56
MAUGIF 480x640 0.00569 1.76 10.63
4.6 Ablation studies
Next, we evaluate the effectiveness of DCIAE framework and $\psi$. Initially, we verify whether DCIAE framework can effectively extract the common features and their respective specific features of the source images. We present the visual results in Figure 9.
Figure 9: Visual demonstration of the fusion process.
For MFF, we find that the extracted common part appears as a fully blurred image. In VIF, the shared components are minimal due to the significant modality gap between the source images. These are consistent with our intuition, which can be primarily attributed to loss2 in (6). Moreover, the modality specific features of the source images highlight their distinct contributions to the fusion process. This verifies that our method is not a ‘black box’, but rather a transparent architecture capable of revealing the contributions of each input modality during fusion.
Next, we validate the effectiveness of the truncation function $\psi$ described in equation (2).
Figure 10: Ablation study on $\psi$ in the VIF task.
Figure 10 illustrates that, without $\psi$, the extracted IR-specific features still contain a significant amount of irrelevant environmental information, which is predominantly composed of negative values. This causes the fused image to appear noticeably darker. When $\psi$ is incorporated, the important regions in the IR image are highlighted while less relevant ones are suppressed, resulting in a fused image with virtually no color cast. This fully validates the effectiveness of $\psi$.
5 Conclusion
In this paper, we propose a mechanism-aware general image fusion method, termed MAUGIF, which is built upon the DCIAE framework. According to the inherent mechanisms of different fusion tasks, we introduce a classification of additive and multiplicative fusion. Then, we employ the DCIAE framework to decompose source images into common components and modality-specific features. During the fusion process, the decoders are regarded as feature injectors to generate the final fusion images. By aligning the decoder architectures with specific fusion strategies, our design not only supports flexible and effective fusion but also enhances interpretability. Finally, extensive numerical experiments further validate the effectiveness of our method.

### v2

@@
 Extensive experiments are conducted on diverse fusion tasks to validate the effectiveness and generalization ability of our method.
-The code is available at https://anonymous.4open.science/r/MAUGIF.
+The code is available at https://github.com/warren-wzw/MAUGIF.git.
 1 Introduction
@@
 The VIF integrates thermal radiation data from infrared sensors with rich texture and color information from visible cameras.
-This results in the fused images that simultaneously highlight salient objects and retain visual fidelity [Ma2021GANMcCAG].
+This results in fused images that simultaneously highlight salient objects and retain visual fidelity [Ma2021GANMcCAG].
 The MFF task aims at generating an all-in-focus image from a series of partially focused images [Quan2025].
@@
 Task-specific methods are specially designed for a particular type of image fusion task.
-These methods typically leverage prior image information such as sparsity [OTD] and structural sparsity [S4], or deep image features [ZSL] to enhance the fusion performance [Zhang2023].
+These methods typically leverage prior image information such as sparsity [OTD] and structural sparsity [S4], or deep image features [ZSL] to enhance fusion performance [Zhang2023].
 However, these methods often exhibit limited generalization capability across different fusion tasks.
@@
 These methods generally use general features including frequency coefficients [Wang2024a] and deep features [Liang2024FusionFD] to optimize the model.
-Nevertheless, most existing methods fail to fully account for the underlying mechanisms of different tasks, thereby limiting their adaptability and interpretability [Zhu2024TaskCustomizedMO].
+Nevertheless, most existing methods fail to fully account for the underlying mechanisms of different tasks, limiting their adaptability and interpretability [Zhu2024TaskCustomizedMO].
 To address these issues, we propose a mechanism-aware unsupervised general image fusion (MAUGIF) method based on Dual Cross-Image Autoencoders (DCIAE).
@@
 By proposing a classification of fusion into additive and multiplicative types, we can gain insight into how the modality-specific features are integrated.
-This not only guides the architectural design of the network, but also enhances model generalization and interpretability.
+This not only guides the architectural design of the network, but also enhances the model’s generalization and interpretability.
   • Comprehensive validation across diverse fusion tasks.
 Extensive numerical experiments are conducted on various fusion tasks (HMF, VIF, MFF and MEF) to validate the effectiveness and generalization capability of the proposed general image fusion method.
-Figure 2: Visual illustration of the fusion process for MFF.
+Figure 2: Visual illustration of the fusion process for MFF task.
 Figure 3: The mechanism-aware DCIAE framework for general image fusion, illustrated with four scenarios.
@@
 2.1 Task-specific image fusion methods
-Task-specific fusion methods are designed for a particular fusion task, leveraging the inherent properties of image and priors to enhance performance.
-Model-based fusion methods typically rely on mathematical modeling and optimization to exploit the priors of images [NSSR, OTD].
+Task-specific fusion methods are designed for a particular fusion task, leveraging the inherent properties of input images and prior knowledge to enhance performance.
+Model-based fusion methods typically rely on mathematical modeling and optimization to exploit image priors [NSSR, OTD].
 For example, these methods often leverage various matrix or tensor decompositions to capture and exploit the intrinsic structure of image data [CP1, CSTF].
-In addition, they can incorporate image priors via designed regularizers, such as low-rankness [Hysure], sparsity [WANG2014477], smoothness [Ma2016], and non-local self-similarity [LTMR].
-Deep learning-based methods capture high-level semantic content by learning hierarchical representations directly from data.
-Numerous methods employ convolutional neural networks (CNNs) to extract such features from source images [ZSL, Li2019].
+In addition, they can incorporate image priors using handcrafted regularizers, such as low-rankness [Hysure], sparsity [WANG2014477], smoothness [Ma2016], and nonlocal self-similarity [LTMR].
+Deep learning-based methods capture high-level semantic content by learning hierarchical representations directly from images.
+Numerous methods employ convolutional neural networks (CNN) to extract such features from source images [ZSL, Li2019].
 Subsequently, generative adversarial networks (GANs) [Ma2019FusionGANAG, Zhang2021a] and vision Transformers [Shao2024, Liu2024] have been introduced to better preserve structural details and global contextual information.
-More recently, diffusion-based fusion methods have also attracted significant attention [Yue2023, Wu2025].
+Diffusion-based fusion methods have also attracted significant attention [Yue2023, Wu2025].
 Furthermore, high-level vision tasks including object detection and semantic segmentation, have been incorporated to guide the fusion process [Tang2022ImageFI, Wang2025DiFusionSegDS].
-These methods often deliver optimal performance on specific fusion tasks but exhibit limited generalizability.
+These methods often deliver high performance on specific fusion tasks but exhibit limited generalizability.
 2.2 General image fusion methods
@@
 Recently, Zhang et al.
-[Zhang2020a] proposed a supervised training framework for MFF, which can be generalized to other tasks by adjusting fusion conditions.
+[Zhang2020a] proposed a supervised training framework for MFF task, which can be generalized to other tasks by adjusting fusion conditions.
 Ma et al.
@@
 To capture the diverse fusion mechanisms of information integration, we classify image fusion into additive and multiplicative types.
-The classification depends on whether the injected modality-specific features are added to the target or modulate it.
+The classification depends on whether modality-specific features are added to the target or modulate it.
 The specific definitions are as follows.
@@
 $\vskip-2.84544pt\textbf{F}=\textbf{Y}+\psi(\Delta_{x})\quad\text{or}\quad\textbf{F}=\textbf{X}+\psi(\Delta_{y}),$ (1)
-where $\psi$ is a nonexpansive projection, i.e., $\|\psi(x)-\psi(y)\|\leq\|x-y\|$, $\Delta_{x}:=\textbf{X}-\textbf{C}$ and $\Delta_{y}:=\textbf{Y}-\textbf{C}$ are the modality-specific features unique to X and Y, respectively, and C is the common content of the same size as X and Y.
+where $\psi$ is a non-expansive projection, i.e., $\|\psi(x)-\psi(y)\|\leq\|x-y\|$, $\Delta_{x}:=\textbf{X}-\textbf{C}$ and $\Delta_{y}:=\textbf{Y}-\textbf{C}$ are modality-specific features unique to X and Y, respectively, and C is the common content of the same size as X and Y.
 For example, in visible-infrared image fusion, the goal is to highlight the thermal radiation of the target from the infrared image (IR) while preserving the structural content of the visible image, which is typically considered additive fusion.
@@
 (The choice of $\psi$) The design of $\psi$ is adapted to the specific fusion task.
-For MFF where the modality-specific information of images is all needed, $\psi$ can be set to the identity mapping, i.e.
-$\psi(x)=x$, to preserve all fine details.
+For MFF where the modality-specific information of images is all needed, $\psi$ can be set to the identity mapping, i.e., $\psi(x)=x$, to preserve all fine details.
 For VIF, a hard-thresholding:
 $\psi(x)=x\quad\text{if}~~x>\sigma;\quad\text{otherwise}~~0,\vskip-2.84544pt$ (2)
-can be selected to suppress weak or spurious responses and enhance salient thermal targets, where $\sigma\in[0,~0.4]$.
+can be employed to suppress weak or spurious responses and enhance salient thermal targets, where $\sigma\in[0,~0.4]$.
 Definition 3.2
@@
 Based on this, we propose the dual cross-image autoencoder (DCIAE) framework for general image fusion, as displayed in Figure 3. The process can be divided into three steps:
-1) To extract the common components of source images, we employ two encoders, denoted as $E_{x}(\cdot)$ and $E_{y}(\cdot)$, to encode X and Y respectively, such that $E_{x}(\textbf{X})\approx E_{y}(\textbf{Y})$.
+1) To extract the common components of source images, we employ two encoders, denoted as $E_{x}(\cdot)$ and $E_{y}(\cdot)$, to encode X and Y, respectively, such that $E_{x}(\textbf{X})\approx E_{y}(\textbf{Y})$.
 The loss function for the encoders can be formulated as:
@@
 This aims primarily to extract common content that aligns with visual intuition.
-2) Two decoders, denoted as $D_{x}(\cdot)$ and $D_{y}(\cdot)$, are employed to decode the learned shared content.
+2) Two decoders, denoted by $D_{x}(\cdot)$ and $D_{y}(\cdot)$, are employed to decode the learned shared content.
 The objective is to reconstruct the source images X and Y from the common components, i.e.,
 $D_{x}(E_{x}(\textbf{X}))=\textbf{X},\quad D_{y}(E_{y}(\textbf{Y}))=\textbf{Y}.$ (7)
-The decoder acts as a ‘feature injector’, which injects the source image’s modality-specific features into the shared content, thereby reconstructing the source images.
+The decoder acts as a feature injector, which injects the source image’s modality-specific features into the shared content, thereby reconstructing the source images.
 The loss function for the decoders can be defined as:
@@
 Figure 4: Architecture of the Encoder and Decoder.
-3) Finally, we perform image fusion by injecting the modality-specific feature of image X into image Y, or vice versa. Therefore, the fused image F can be formulated as:
+3) Finally, we perform image fusion by injecting the modality-specific features of image X into image Y, or vice versa. Therefore, the fused image F can be formulated as:
 $\textbf{F}=D_{x}(\textbf{Y})\quad\text{or}\quad\textbf{F}=D_{y}(\textbf{X}).$ (9)
-In this process, the decoder serves directly as a fusion operator, shifting the fusion process from conventional rule-based design to intrinsic feature injection.
+In this process, the decoder serves directly as a fusion operator, shifting the fusion process from conventional rule-based design to the injection of modality specific feature.
 Additive fusion and multiplicative fusion correspond to different decoder architectures, as shown in Figure 3 (e).
-In additive fusion, the decoder architecture follows a residual paradigm, where the modality-specific features of X are injected as an additive manner, i.e.,
+In additive fusion, the decoder architecture follows a residual paradigm, where the modality-specific feature of X is injected as an additive manner, i.e.,
 $\vskip-1.42271ptD_{x}(\cdot)=(\cdot)+\psi(\Delta_{x}).$ (10)
@@
 Remark 3.2
-Under DCIAE framework, image fusion is no longer a ‘black-box’ process.
+In DCIAE framework, image fusion is no longer a ‘black-box’ process.
 We can clearly observe the modality-specific features contributed by the source images during the fusion process.
@@
 Next, we analyze the factors affecting DCIAE’s performance. Taking additive fusion as an example, we denote
-  • $\mathcal{C}^{*}$: the ground-truth shared content between X and Y.
+  • $\mathcal{C}^{*}$: the ground truth shared content between X and Y.
   • $\widehat{\textbf{F}}$: the fusion result of DCIAE under ideal decoder, i.e., $D_{x}(E_{x}(\textbf{X}))=\textbf{X}$ holds exactly.
@@
 4.1.1 Implementation details
-All experiments are conducted on Ubuntu 20.04 with the PyTorch 2.2.0.
+All experiments are conducted on Ubuntu 20.04 with PyTorch 2.2.0.
 The hardware configuration includes an Intel Core i7-13700KF CPU with a clock speed of 3.4GHz, 32GB of RAM, and an NVIDIA GeForce RTX 4090 GPU with 24GB of VRAM.
@@
 CAVE1 and Harvard Dataset [data11] for HMF: For the CAVE dataset, we selected ‘Toys’ and ‘Peppers’, whose spatial dimensions were downsampled by different scale factors (sf).
-We simulated the MSIs with three bands by the Nikon D700 camera2.
-For Harvard dataset, the ‘img3’ and ‘imgc6’ images are selected.
+We simulated MSIs with three bands by the Nikon D700 camera2.
+For the Harvard dataset, images ‘img3’ and ‘imgc6’ are selected.
 Each contains a 31-band HSI, which serves as the ground truth.
-Then, the low-resolution HSIs are generated through downsampling the original HSI at different scale factors.
+The low-resolution HSIs are generated through downsampling the original HSI at different scale factors.
 We divide the spectral bands of HSI into three parts and generate a 3-band MSI by averaging the bands.
-We corrupt the HSI and MSI with Gaussian noise to obtain SNRs of 35 dB and 40 dB, respectively.
-MSRS Dataset [Tang2022ImageFI] for VIF: This dataset is a benchmark for infrared-visible image fusion and segmentation.
+We corrupt HSI and MSI with Gaussian noise to obtain SNRs of 35 dB and 40 dB, respectively.
+MSRS Dataset [Tang2022ImageFI] for VIF: This dataset is a benchmark for the fusion and segmentation of infrared-visible images.
 It comprises 1,444 well-registered image pairs.
@@
 Harvard Dataset3 for MEF: This dataset is a well-known medical imaging benchmark for multimodal fusion.
-It provides a large number of CT–MRI, PET–MRI, and SPECT–MRI image pairs, which are widely used for training and evaluating medical image fusion models.
+It provides a large number of CT–MRI, PET–MRI, and SPECT–MRI image pairs, which are widely used to train and evaluate medical image fusion models.
 11footnotetext: https://www.cs.columbia.edu/CAVE/databases/multispectral/ 22footnotetext: https://maxmax.com/spectral_response.htm33footnotetext: http://www.med.harvard.eduAANLIB/home.html
@@
 Therefore, we select five blind HMF methods for comparison, including the classical CNMF [CNMF] and HySure [Hysure], as well as the recently proposed ZSL [ZSL], TriDLR [Yang2025], and DTDNML [Wang2024].
-Table 1: The evaluation metrics of the fused images for HMF.
+Table 1: The evaluation metrics of the fused images for HMF task.
 Dataset Method PSNR SSIM ERGAS SAM Time(s)
@@
 MAUGIF 40.76 0.969 0.853 5.03 10.8
-Table 1 presents the fusion performance of the compared methods on four HSI datasets.
-From the table, the CNMF and HySure exhibit relatively weaker performance, particularly under large scale factor cases.
-In this cases, the fusion model becomes severely ill-posed, which limits their performance.
+Table 1 presents the fusion performance of the compared methods in four HSI datasets.
+From the table, the CNMF and HySure exhibit relatively weaker performance, particularly in large scale factor cases.
+In these cases, the fusion model becomes severely ill-posed, which limits their performance.
 TriDLR reduces the number of parameters through tensor triple decomposition.
@@
 In contrast, the proposed MAUGIF method exhibits greater robustness to large sf, as the proposed method can perform spectral super-resolution on the MSI, which is less sensitive to spatial downsampling.
-Overall, MAUGIF achieves the optimal performance in four datasets, which demonstrates its effectiveness in HMF.
+Overall, MAUGIF achieves optimal performance in four datasets, which demonstrates its effectiveness in HMF task.
 Figure 7: Experimental results of compared methods on three multi-focus image datasets and the corresponding PSNR values.
@@
 Figure 8: Experimental results of compared methods on three types of medical image datasets. (a)–(c) correspond to different modalities.
-Figure 5 presents the fused images and error maps on the ‘Toys’ dataset for visual comparison.
+Figure 5 presents the fused images and error maps in the ‘Toys’ dataset for visual comparison.
 As shown in the figure, the fused images of HySure exhibit noticeable color distortions, which indicates a significant estimation error in the spectral degradation operator.
@@
 4.3 Comparison on VIF task
-To evaluate the performance of the MAUGIF on visible-infrared image fusion task, we conduct experiments on MSRS [Tang2022ImageFI] dataset, and compare the MAUGIF with several state-of-the-art (SOTA) fusion methods.
-The quantitative results presented in Figure 2 demonstrate that MAUGIF achieves five top rankings and one second-place ranking among all competing methods.
+To evaluate the performance of MAUGIF on visible-infrared image fusion task, we conduct experiments on MSRS [Tang2022ImageFI] dataset, and compare the MAUGIF with several state-of-the-art (SOTA) fusion methods.
+The quantitative results presented in Table 2 demonstrate that MAUGIF achieves five top rankings and one second-place ranking among all competing methods.
 Specifically, it records an MI of 5.638, an SCD of 1.731, a VIF of 0.949, a $Q^{AB/F}$ score of 0.652, and an SSIM of 0.482.
 These results indicate that MAUGIF effectively preserves mutual information, structural similarity, and visual information fidelity, suggesting superior retention of texture, structure, and perceptual quality.
-Table 2: Quantitative comparisons on image pairs from the MFI-WHU and Harvard datasets.
+Table 2: Quantitative comparisons of six metrics on 361 image pairs from the MSRS dataset.
 Category MI↑ AG↑ SCD↑ VIF↑ $Q^{AB/F}$ SSIM↑
@@
 In summary, both quantitative and qualitative results demonstrate that MAUGIF preserves complementary information from the source images while maintaining visual fidelity.
-The fused images retain key details from both visible and infrared modalities, ensuring high visual quality and enhancing the overall integrity of the fusion process.
 4.4 The performance on MFF and MEF
@@
 This indicates that the fused image preserves both mutual information and visual information fidelity, which fully validates the effectiveness and generalization of our approach.
-Table 3: Quantitative comparisons of six metrics on 361 image pairs from the MSRS dataset.
+Table 3: Quantitative comparisons on image pairs from the MFI-WHU and Harvard datasets.
 Method U2Fusion20 SDNet21 TC-MOA24 MAUGIF
@@
 In VIF, the shared components are minimal due to the significant modality gap between the source images.
-These are consistent with our intuition, which can be primarily attributed to loss2 in (6).
-Moreover, the modality specific features of the source images highlight their distinct contributions to the fusion process.
+These are consistent with our intuition, which can be primarily attributed to $loss_{2}$ in (6).
+Moreover, the modality specific features of source images highlight their distinct contributions to the fusion process.
 This verifies that our method is not a ‘black box’, but rather a transparent architecture capable of revealing the contributions of each input modality during fusion.